OpenStudy (anonymous):

simplify this expression. (4jk^4)^3(jk)^4

OpenStudy (agent0smith):

\[\Large (4jk^4)^3(jk)^4 = 4^3j^3k^{12} \times j^3 k^4\]But you still have to finish multiplying it all together and simplify.

OpenStudy (anonymous):

:p.